A student scored 483 points on a quiz and the students standard or (z) score was -1. the standard deviation of quiz scores was 12. What was the mean of the quiz score?
Answer by stanbon(75887)

A student scored 483 points on a quiz and the students standard or (z) score was -1. the standard deviation of quiz scores was 12. What was the mean of the quiz score?
-----
x = z*s + u
---
u = x - z*s
u = 483-(-1*12)
u = 495
==============
Cheers,
Stan H.
--------------